# From Fog to Failure: How Dehazing Can Harm Clear Image Object Detection

### v2

@@ Abstract
This study explores the challenges of integrating human visual cue-based dehazing into object detection, given the selective nature of human perception. While human vision adapts dynamically to environmental conditions, computational dehazing does not always enhance detection uniformly. We propose a multi-stage framework where a lightweight detector identifies regions of interest (RoIs), which are then enhanced via spatial attention-based dehazing before final detection by a heavier model. Though effective in foggy conditions, this approach unexpectedly degrades the performance on clear images. We analyze this phenomenon, investigate possible causes, and offer insights for designing hybrid pipelines that balance enhancement and detection. Our findings highlight the need for selective preprocessing and challenge assumptions about universal benefits from cascading transformations. The implementation of the framework is available111https://github.com/ashu1069/perceptual-piercing.
@@ 1 Introduction
Low-visibility conditions, such as rain, snow, fog, smoke, and haze, pose significant challenges for deep learning applications in autonomous vehicles, security and surveillance, maritime navigation, and agricultural robotics. Under these conditions, object detection models struggle due to reduced contrast and obscured features, leading to performance degradation. This study proposes a deep learning framework inspired by human visual perception to enhance object recognition in adverse visibility scenarios, particularly in foggy environments.
A key motivation for this work comes from the impact of poor visibility on airport operations, where disruptions in taxiing and docking cause delays and increase reliance on ground support. Limited staff availability exacerbates these issues, sometimes resulting in flight cancellations. While initially driven by airport challenges, the proposed framework applies broadly to various low-visibility scenarios across multiple domains.
@@ 2 Related Work
Despite these advances, challenges persist, including computational complexity (Zhang et al., 2018; Atom et al., 2020; Tang et al., 2022), performance issues under extreme conditions (Khattak et al., 2019; Boban et al., 2012), overfitting due to limited datasets (Zhang et al., 2018; Khattak et al., 2019), and insufficient real-world validation (Liu et al., 2014; Boban et al., 2012; Tang et al., 2022). Some works lack clarity in explanation or rigorous validation (Kramer et al., 2014; Zhan et al., 2023).
@@ 5 Discussions
The unexpected performance improvement of AOD-NetX-based models under foggy conditions (Table 1) raises critical questions about feature adaptation in hybrid object detection pipelines. Typically, object detectors exhibit a decline in accuracy when transitioning from clear to foggy conditions, as seen in conventional YOLOv5x and YOLOv8x models. However, our proposed AOD-NetX integration leads to an inverse trend—where object detection improves in foggy conditions relative to clear ones. This suggests that AOD-NetX introduces an implicit domain adaptation effect, making the detection network more attuned to foggy environments at the cost of generalization to clear images.
One possible explanation lies in dataset bias and overprocessing. Since AOD-NetX is trained primarily on foggy images, its learned feature space is optimized for haze removal but lacks the necessary constraints to preserve features in clear conditions. Consequently, when applied to clear images, the model introduces distortions instead of enhancements, disrupting feature consistency for the object detector. This emphasizes the need for context-aware enhancement, where image processing techniques are selectively applied based on scene conditions rather than indiscriminately.
Furthermore, the results challenge the assumption that cascading pipelines—where lightweight detection informs region-specific enhancement before a final robust detection—always improve performance. While effective in foggy settings, this multi-stage approach appears to introduce trade-offs, potentially harming accuracy in clear conditions. Future designs must strike a balance between specialization for adverse weather conditions and adaptability to diverse environments. Another consideration is the real-time feasibility of this approach. RoI-specific dehazing adds computational overhead, which could limit deployment in time-sensitive applications such as autonomous driving. Optimizing processing efficiency while retaining performance gains remains an open challenge.

### v4

# From Fog to Failure: The Unintended Consequences of Dehazing on Object Detection in Clear Images

@@ Abstract
This study explores the challenges of integrating human visual cue-based dehazing into object detection, given the selective nature of human perception. While human vision adapts dynamically to environmental conditions, computational dehazing does not always enhance detection uniformly. We propose a multi-stage framework where a lightweight detector identifies regions of interest (RoIs), which are then improved via spatial attention-based dehazing before final detection by a heavier model. Though effective in foggy conditions, this approach unexpectedly degrades the performance on clear images. We analyze this phenomenon, investigate possible causes, and offer insights for designing hybrid pipelines that balance enhancement and detection. Our findings highlight the need for selective preprocessing and challenge assumptions about universal benefits from cascading transformations. The implementation of the framework is available here111https://github.com/ashu1069/perceptual-piercing.
@@ 1 Introduction
Low-visibility conditions, such as rain, snow, fog, smoke, and haze, pose significant challenges for deep learning applications in autonomous vehicles, security and surveillance, maritime navigation, and agricultural robotics. Object detection models struggle in these environments due to reduced contrast and obscured features, often leading to performance degradation. While image enhancement methods, including dehazing, improve visibility, they can also introduce artifacts or distortions that negatively impact downstream tasks. Overprocessing may lead to false positives and increased computational overhead, highlighting the need for more selective enhancement strategies. Motivated by real-world challenges, such as disruptions in airport operations where poor visibility delays taxiing and docking, this study proposes a vision-inspired deep learning framework tailored for adverse conditions, particularly fog.
@@ 2 Related Work
Despite these advances, challenges persist, including computational complexity (Zhang et al., 2018; Atom et al., 2020; Tang et al., 2022), performance issues under extreme conditions (Khattak et al., 2019; Boban et al., 2012), overfitting due to limited datasets (Zhang et al., 2018; Khattak et al., 2019), and insufficient real-world validation (Liu et al., 2014; Boban et al., 2012; Tang et al., 2022). Some works lack rigorous validation (Kramer et al., 2014; Zhan et al., 2023).
@@ 5 Discussions
The unexpected performance improvement of AOD-NetX-based models under foggy conditions (Table 1) raises critical questions about feature adaptation in hybrid object detection pipelines. Typically, object detectors show a decrease in accuracy when transitioning from clear to foggy conditions, as seen in conventional YOLOv5x and YOLOv8x models. However, our proposed AOD-NetX integration leads to an inverse trend, where object detection improves under foggy conditions relative to that under clear conditions. This suggests that AOD-NetX introduces an implicit domain adaptation effect, which makes the detection network more attuned to foggy environments at the cost of generalization to clear images.
One possible explanation lies in bias and overprocessing in the data set. Since AOD-NetX is trained primarily on foggy images, its learned feature space is optimized for haze removal, but lacks the necessary constraints to preserve features in clear conditions. Consequently, when applied to clear images, the model introduces distortions instead of enhancements, disrupting feature consistency for the object detector. This emphasizes the need for context-aware enhancement, where image-processing techniques are selectively applied based on scene conditions rather than indiscriminately.
Furthermore, the results challenge the assumption that cascading pipelines, where lightweight detection informs region-specific enhancement before a final robust detection, always improve performance. While effective in foggy settings, this multi-stage approach appears to introduce trade-offs, potentially harming accuracy in clear conditions. Future designs must strike a balance between specialization for adverse weather conditions and adaptability to diverse environments. Another consideration is the real-time feasibility of this approach. RoI-specific dehazing adds computational overhead, which could limit deployment in time-sensitive applications such as autonomous driving. Optimizing processing efficiency while retaining performance gains remains an open challenge.